# A Lanczos-Based Algorithmic Approach for Spike Detection in Large Sample Covariance Matrices

### v1

@@ 6. Statistical consistency of our proposed estimator
Suppose that the sample covariance matrix $W$, as given in (1.1), satisfies Assumption 2. With $k=1$, suppose Lanczos is run for $n=\lceil C\log N\rceil$ steps for $C>0$ sufficiently large. Let $\widehat{m}_{0}(z)$ be the output of Estimation Algorithm P.2 and fix $20<\delta<1/$ 2. Let $\widehat{r}$ be the number of poles of $\widehat{m}_{0}(z)$ for $z>\widehat{\gamma}_{+}+N^{-\delta}$ and let $\widehat{\gamma}_{j}$, $j=1,2,\ldots,\widehat{r}$ be their locations. Then for every $D>0$ there exists $N_{0}(D)$ such that for $N\geq N_{0}(D)$,
@@
Next, suppose one of the off-diagonal entries of $\mathcal{L}_{0}\mathcal{L}_{0}^{*}$ vanishes. This implies that Lanczos failed to run to completion, implying that either $W$ has a repeated eigenvalue, or one of the projections of $\mathbf{b}_{1}$ in Estimation Algorithm P.2 onto one of the eigenvectors of $W$ vanishes — i.e. the VESD $\mu_{W,\mathbf{b}}$ is degenerate. Using [9] along with [46], for example, this degeneracy occurs with probability $\mathrm{O}(N^{-D})$ for any $D>0$ — so that with overwhelming probability $\mathcal{L}_{0}\mathcal{L}_{0}^{*}$ has non-zero off-diagonal entries. From the theory of Jacobi operators [10], with overwhelming probability, the discrete eigenvalues of $\mathcal{L}_{0}\mathcal{L}_{0}^{*}$ are simple and $\widehat{m}_{0}(z)$ will have a pole at every discrete eigenvalue of $\mathcal{L}_{0}\mathcal{L}_{0}^{*}$. So, $\widehat{m}_{0}(z)$ will have a pole within a distance $\tau$ of an asymptotic outlier $f(-\tilde{\sigma}_{j}^{-1})$. On the event where the ranks of the projectors and are equal and $\mathcal{L}_{0}\mathcal{L}_{0}^{*}$ has non-zero off-diagonal entries, the only way the multiplicity of eigenvalues within a circle of radius $\tau$, centered at $f(-\tilde{\sigma}_{j}^{-1})$, would be miscounted by counting the poles of $\widehat{m}_{0}(z)$ is if $\mathcal{L}_{0}\mathcal{L}_{0}^{*}$ had an eigenvalue of multiplicity greater than one. This has been ruled out. ∎
@@ 7. Numerical Experiments and Comparisons
Throughout this section, we examine our approach using $C=1$ and $\delta=0.25$ in Algorithm P.3, as they yield reliable results. Based on several experiments and the fact that our support estimates exhibit fluctuations of order $N^{-1/2}$ , we find that values of $\delta$ closer to $1/2$ tend to overestimate the number of spikes, while values near 00 tend to underestimate. A more detailed analysis of the choice of parameters is left for future work.
We also compare our approach to other methods, including BEMA0, BEMA [29], DDPA [20], and the eigen-gap method of Passemier and Yao (Pass&Yao) [40]. The BEMA0 method constructs the right endpoint using bulk eigenvalues and applies a correction derived from the distribution of the largest eigenvalue. BEMA approximates the non-spiked covariance matrix $\Sigma_{0}$ as a diagonal matrix with iid entries from a Gamma distribution. It then uses Monte Carlo simulations to estimate the distribution of the largest eigenvalue, which is used to build a threshold for the spikes. DDPA, on the other hand, relies on parallel analysis, estimating the non-spiked covariance matrix using the diagonal of the sample covariance and establishing a threshold through a deterministic recursive procedure. Finally, the Pass&Yao method approximates $\Sigma_{0}$ by $\sigma^{2}I$, where $\sigma^{2}$ is estimated using maximum likelihood. The number of spikes is then identified by comparing the spectral gaps to a threshold derived from Monte Carlo simulations.
@@
Estimated outliers Error $5.26704$ $141.42108e\!-\!$ 14 $5.24308$ $152.66453e\!-\!$ 15 $4.72643$ $155.32907e\!-\!$ 15
Estimated outliers Error $6.05801$ $121.59250e\!-\!$ 12 $6.01288$ $131.10134e\!-\!$ 13 $5.66482$ $134.32542e\!-\!$ 13
Estimated outliers Error $7.04843$ $123.89999e\!-\!$ 12 $6.79924$ $112.50874e\!-\!$ 11 $6.38327$ $93.21105e\!-\!$ 9
Figure 3. The rows correspond to $c=0.1$, $c=0.5$, and $c=0.9$, respectively. Left: Estimated outliers and ASD obtained using Algorithms P.2 and P.3 with $k=200$ vectors. These estimates are compared to the ESD of $W$ from Simulation 1 and the MP density given in (7.3). Right: Estimated outlier locations and the corresponding error between the estimates and the true outliers of $W$.
The behavior of $W$ is analyzed for three values of $c$: $0.1$, $0.5$, and $0.9$, representing different regimes within the range $c\in(0,1)$. For each case, the estimated ASD and outliers of $W$ are visualized with $N=5000$, and the number of vectors is chosen as $k=100$ in Algorithms P.2 and P.3. Figure 3 shows that the approximate ASD closely matches the exact density, and the detected outliers align well with the true ones. The error in estimating the spike locations is of the order $10^{-14}$ and $10^{-12}$ for $c=0.1$ and $c=0.5$, respectively. However, when the spikes approach $\gamma_{+}$, as in the case of $c=0.9$, the accuracy decreases.
Next, we investigate the accuracy of Algorithm P.2 in estimating both the density endpoints and the density itself, with the number of vectors $k$ fixed at 100. This time, we consider the three scenarios for $c$ while varying $N$ from $100$ to $8000$. For each value of $N$, the errors are averaged over 200 trials. In each trial, the error in the support is calculated as $\max\{|\gamma_{+}-\widehat{\gamma}_{+}|,|\gamma_{-}-\widehat{\gamma}_{-}|\}$, and the density error is determined as
@@
Figure 4. Comparison of the estimated support and density from Algorithm P.2 with $k=100$ against the true counterparts from the MP law. The plots illustrate the convergence of errors in Simulation 1 for $c=0.1$, $0.5$, and $0.9$ as $N$ increases. Errors are averaged over 200 trials, with each point representing the mean error and vertical error bars indicating the standard deviation.
Finally, we examine the accuracy of our method for spike detection. To highlight its effectiveness, we consider one vector ($k=1$) in Algorithm P.3. As before, we consider the three regimes $c=0.1,0.5,0.9$ and vary $N$. For each value of $N$, we generate 200 samples of the covariance matrix $W$ and evaluate both the probability of correctly identifying the number of spikes and the average number of detected spikes. We also demonstrate the efficiency of our approach by comparing its average runtime with the time required to compute eigenvalues across the 200 samples. The results are presented in Figure 5 and Table 1. Our estimator for $\widehat{r}$ exhibits greater accuracy for $c=0.1$ and $c=0.5$ than for $c=0.9$, which can be attributed to the larger gap between the spikes and the density endpoints. Nonetheless, the accuracy improves as $N$ increases and eventually reaches 1 in all cases. Furthermore, even when the probability of correctly estimating the number of spikes is low, the average value of $\widehat{r}$ remains relatively close to the true spike count. These results also highlight the efficiency of the detection algorithm, with its runtime remaining small as $N$ grows, in contrast to the substantial increase in the computational cost of eigenvalue computation.
Figure 5. Left: Spike detection accuracy in Simulation 1 for $c=0.1$, $0.5$, and $0.9$ as $N$ varies, based on 200 sample realizations with the algorithm run using $k=1$. Right: Comparison of the average computation time for eigenvalue calculation and spike detection across the samples for $c=0.5$.
$N$ $c=0.1$ $c=0.5$ $c=0.9$
$200$ $3.04$ $(0.97)$ $4.04$ $(0.45)$ $7.05$ $(0.17)$
$2000$ $3.00$ $(1.00)$ $3.06$ $(0.95)$ $3.35$ $(0.81)$
$4000$ $3.00$ $(1.00)$ $3.01$ $(0.99)$ $3.17$ $(0.89)$
$6000$ $3.00$ $(1.00)$ $3.00$ $(0.99)$ $3.13$ $(0.91)$
$8000$ $3.00$ $(1.00)$ $3.00$ $(1.00)$ $3.08$ $(0.95)$
Table 1. Estimated number of spikes for $W$ in Simulation 1 for different values of $c$ and $N$, using a single vector ($k=1$) in the detection algorithm. Each table entry represents the average over 200 samples, with the value in brackets denoting the probability of correctly detecting the true number of spikes.
@@
We set $N=8000$ and $M=16000$, with $X\in\mathbb{R}^{N\times M}$ consisting of iid standard normal entries  [^7] Similar results were obtained when the entries of $X$ followed a Rademacher distribution or a Beta distribution with shape parameters $\alpha=\beta=1/2$.. The parameter $\delta$ ranges from $1.5$ to $3$, crossing the BBP transition, where the number of spikes increases from $2$ to $3$ at $\delta=1+\sqrt{c}$ with $c=N/M$. We evaluate the performance of Algorithm P.3 with $k=1,50,100,200$, alongside BEMA0 and DDPA, in estimating the number of spikes across $100$ realizations of the sample covariance matrix. In addition, we examine the average number of detected spikes and compare the computational efficiency of each method based on their average runtime. The results are summarized in Figures 6, 7 and Table 2.
Figure 6. Spike detection accuracy in Simulation 2, averaged over 100 trials for each value of $\delta$ as it varies from 1 to 3. During this process, the sample covariance matrix undergoes the BBP transition, increasing the number of spikes from 2 to 3. We compare the performance of Algorithm P.3 with different numbers of vectors against the BEMA0 and DDPA methods.
Figure 7. Comparison of the average computation time in Simulation 2, measured over 100 trials, for Algorithm P.3 with varying numbers of vectors, alongside BEMA0 and DDPA. The runtimes of BEMA0 and DDPA are nearly identical, as both are primarily determined by the cost of eigenvalue computation.
@@
$\delta$ Lanczos . with . $k=1$ Lanczos . with . $k=50$ Lanczos with $k=100$ Lanczos with $k=200$ BEMA0 DDPA
$1.50$ $2.00$ $(1.00)$ $2.00$ $(1.00)$ $2.00$ $(1.00)$ $2.00$ $(1.00)$ $2.16$ $(0.84)$ $2.00$ $(1.00)$
$1.75$ $2.00$ $(0.00)$ $2.00$ $(0.00)$ $2.00$ $(0.00)$ $2.00$ $(0.00)$ $2.61$ $(0.59)$ $2.00$ $(0.00)$
$1.90$ $2.01$ $(0.01)$ $2.01$ $(0.01)$ $2.00$ $(0.00)$ $2.02$ $(0.02)$ $3.07$ $(0.93)$ $2.00$ $(0.00)$
$2.00$ $2.21$ $(0.21)$ $2.29$ $(0.29)$ $2.21$ $(0.21)$ $2.23$ $(0.23)$ $3.08$ $(0.92)$ $2.00$ $(0.00)$
$2.10$ $2.92$ $(0.92)$ $2.84$ $(0.84)$ $2.84$ $(0.84)$ $2.93$ $(0.93)$ $3.08$ $(0.92)$ $2.00$ $(0.00)$
$2.25$ $3.00$ $(1.00)$ $3.00$ $(1.00)$ $3.00$ $(1.00)$ $3.00$ $(1.00)$ $3.08$ $(0.92)$ $2.00$ $(0.00)$
$2.50$ $3.00$ $(1.00)$ $3.00$ $(1.00)$ $3.00$ $(1.00)$ $3.00$ $(1.00)$ $3.03$ $(0.97)$ $2.00$ $(0.00)$
$2.75$ $3.00$ $(1.00)$ $3.00$ $(1.00)$ $3.00$ $(1.00)$ $3.00$ $(1.00)$ $3.07$ $(0.93)$ $3.00$ $(1.00)$
Table 2. Estimated number of spikes for $W$ in Simulation 2 as $\delta$ crosses the BBP transition, increasing the number of spikes from $2$ to $3$. We compare Algorithm P.3 with different numbers of vectors against BEMA0 and DDPA. Each table entry represents the average over 100 samples, with the value in brackets indicating the probability of correctly identifying the true number of spikes.
@@
Figure 8. Left: The ESD of $\Sigma$, where $\Sigma$ is an $N\times N$ diagonal matrix with $N=2500$ and entries given by the quantiles of the density defined in (7.5). The first two diagonal entries are modified to 7 and 5 forming the spikes of $\Sigma$. Right: The ESD of $W$ defined in Simulation 3. The ESD is compared against the estimated outliers, their locations, and the approximate ASD obtained using Algorithms P.2 and P.3 with $k=200$.
We consider the sample covariance matrix $W$ under the spiked covariance model in (1.1), where $N=2500$, $M=25000$ and $\sqrt{M}X\in\mathbb{R}^{N\times M}$ has iid standard Gaussian entries. The covariance matrix $\Sigma\in\mathbb{R}^{N\times N}$ is diagonal, and its eigenvalues follow a deterministic distribution consisting of a bulk component and additional spikes. The bulk density is expressed as
@@
where $K$ is the normalization constant, and the spikes of $\Sigma$ are located at $7,\delta$. See Figure 8 for a visualization of the ESD of $\Sigma$ and the ASD of $W$ when $\delta=5$. We approximate the ASD and detect the outliers of $W$ using Algorithms P.2 and P.3 with $k=100$ vectors. Our ASD estimate closely matches the empirical distribution, accurately identifying both the location and number of spikes.
Next, we vary the parameter $\delta=5,7,9,21$ and evaluate the performance of our detection algorithm with a single vector ($k=1$), comparing it to other spike detection methods. For each value of $\delta$, we perform the evaluation over 100 samples. Specifically, we examine the accuracy, the average number of spikes detected, and the runtime for each method. The results are summarized in Table 3.
Our algorithm outperforms the other methods, as it consistently captures the correct number of spikes with a probability close to 1. DDPA performs second best, with its performance improving as $\delta$ increases, reaching a probability of 1 when $\delta=21$. This demonstrates the versatility of our approach, as it can handle a wide range of covariance matrices without assuming any specific distribution or structure for $\Sigma$. Additionally, we observe that Algorithm P.3 achieves the shortest runtime, followed by BEMA0 and DDPA, which have similar runtimes as their costs are mainly dominated by the eigenvalue computations. BEMA and Pass&Yao are more computationally demanding, as they rely on Monte Carlo simulations with large matrices.
@@
Probability $0.00$ $0.00$ $0.52$ $0.00$ $0.93$
Average $461.06$ $0.00$ $2.66$ $1.00$ $1.99$
Time (sec) $3.443$ $1993.270$ $1164.536$ $3.444$ $0.088$
@@
Probability $0.00$ $0.00$ $0.48$ $0.70$ $0.99$
Average $461.11$ $0.00$ $2.77$ $1.40$ $2.01$
Time (sec) $3.526$ $2054.871$ $1236.337$ $3.526$ $0.0843$
@@
Probability $0.00$ $0.00$ $0.59$ $1.00$ $1.00$
Average $461.01$ $0.00$ $2.61$ $2.00$ $2.00$
Time (sec) $3.523$ $2044.017$ $1183.847$ $3.523$ $0.070$
@@
Probability $0.00$ $0.00$ $0.60$ $1.00$ $0.98$
Average $461.42$ $1.00$ $2.46$ $2.00$ $2.02$
Time (sec) $3.594$ $2032.415$ $1133.657$ $3.594$ $0.071$
Table 3. Probability of correct estimation, average number of spikes detected, and runtime for different methods in Simulation 3, evaluated over 100 samples for each value of $\delta$.

### v2

@@ 6. Statistical consistency of our proposed estimator
Suppose that the sample covariance matrix $W$, as given in (1.1), satisfies Assumption 2. With $k=1$, suppose Lanczos is run for $n=\lceil C\log N\rceil$ steps for $C>0$ sufficiently large. Let $\widehat{m}_{0}(z)$ be the output of Estimation Algorithm P.2 and fix $0<\delta<1/2$. Let $\widehat{r}$ be the number of poles of $\widehat{m}_{0}(z)$ for $z>\widehat{\gamma}_{+}+N^{-\delta}$ and let $\widehat{\gamma}_{j}$, $j=1,2,\ldots,\widehat{r}$ be their locations. Then for every $D>0$ there exists $N_{0}(D)$ such that for $N\geq N_{0}(D)$,
@@
Next, suppose one of the off-diagonal entries of $\mathcal{L}_{0}\mathcal{L}_{0}^{*}$ vanishes. This implies that Lanczos failed to run to completion, implying that either $W$ has a repeated eigenvalue, or one of the projections of $\mathbf{b}_{1}$ in Estimation Algorithm P.2 onto one of the eigenvectors of $W$ vanishes — i.e. the VESD $\mu_{W,\mathbf{b}}$ is degenerate. Using [Christoffersen2025] along with [Tikhomirov2016], for example, this degeneracy occurs with probability $\mathrm{O}(N^{-D})$ for any $D>0$ — so that with overwhelming probability $\mathcal{L}_{0}\mathcal{L}_{0}^{*}$ has non-zero off-diagonal entries. From the theory of Jacobi operators [DeiftOrthogonalPolynomials], with overwhelming probability, the discrete eigenvalues of $\mathcal{L}_{0}\mathcal{L}_{0}^{*}$ are simple and $\widehat{m}_{0}(z)$ will have a pole at every discrete eigenvalue of $\mathcal{L}_{0}\mathcal{L}_{0}^{*}$. So, $\widehat{m}_{0}(z)$ will have a pole within a distance $\tau$ of an asymptotic outlier $f(-\tilde{\sigma}_{j}^{-1})$. On the event where the ranks of the projectors and are equal and $\mathcal{L}_{0}\mathcal{L}_{0}^{*}$ has non-zero off-diagonal entries, the only way the multiplicity of eigenvalues within a circle of radius $\tau$, centered at $f(-\tilde{\sigma}_{j}^{-1})$, would be miscounted by counting the poles of $\widehat{m}_{0}(z)$ is if $\mathcal{L}_{0}\mathcal{L}_{0}^{*}$ had an eigenvalue of multiplicity greater than one. This has been ruled out. ∎
@@ 7. Numerical Experiments and Comparisons
Throughout this section, we examine our approach using $C=1$ and $\delta=0.25$ in Algorithm P.3, as they yield reliable results. Based on several experiments and the fact that our support estimates exhibit fluctuations of order $N^{-1/2}$ , we find that values of $\delta$ closer to $1/2$ tend to overestimate the number of spikes, while values near 0 tend to underestimate. A more detailed analysis of the choice of parameters is left for future work.
We also compare our approach to other methods, including BEMA0, BEMA [Ke2023], DDPA [PA], and the eigen-gap method of Passemier and Yao (Pass&Yao) [passemier2014estimation]. The BEMA0 method constructs the right endpoint using bulk eigenvalues and applies a correction derived from the distribution of the largest eigenvalue. BEMA approximates the non-spiked covariance matrix $\Sigma_{0}$ as a diagonal matrix with iid entries from a Gamma distribution. It then uses Monte Carlo simulations to estimate the distribution of the largest eigenvalue, which is used to build a threshold for the spikes. DDPA, on the other hand, relies on parallel analysis, estimating the non-spiked covariance matrix using the diagonal of the sample covariance and establishing a threshold through a deterministic recursive procedure. Finally, the Pass&Yao method approximates $\Sigma_{0}$ by $\sigma^{2}I$, where $\sigma^{2}$ is estimated using maximum likelihood. The number of spikes is then identified by comparing the spectral gaps to a threshold derived from Monte Carlo simulations.
@@
Estimated outliers Error $5.25537$ 1.50990e141.50990e\!-\!14 $5.24390$ 1.77635e151.77635e\!-\!15 $4.70768$ 2.04281e142.04281e\!-\!14
Estimated outliers Error $6.09914$ 8.88178e168.88178e\!-\!16 $6.05833$ 1.50990e141.50990e\!-\!14 $5.58658$ 1.24344e141.24344e\!-\!14
Estimated outliers Error $7.23994$ 8.88178e158.88178e\!-\!15 $6.96060$ 3.55271e153.55271e\!-\!15 $6.57313$ 6.21724e156.21724e\!-\!15
Figure 3. The rows correspond to $c=0.1$, $c=0.5$, and $c=0.9$, respectively. Left: Estimated outliers and ASD obtained using Algorithms P.2 and P.3 with $k=100$ vectors. These estimates are compared to the ESD of $W$ from Simulation 1 and the MP density given in (7.3). Right: Estimated outlier locations and the corresponding error between the estimates and the true outliers of $W$.
The behavior of $W$ is analyzed for three values of $c$: $0.1$, $0.5$, and $0.9$, representing different regimes within the range $c\in(0,1)$. For each case, the estimated ASD and outliers of $W$ are visualized with $N=5000$, and the number of vectors is chosen as $k=100$ in Algorithms P.2 and P.3. Figure 3 shows that the approximate ASD closely matches the exact density, and the detected outliers align well with the true ones, with errors on the order of $10^{-14}$.
[…]
@@
Figure 4. Comparison of the estimated support and density from Algorithm P.2 with $k=100$ against the true counterparts from the MP law. The plots illustrate the convergence of errors in Simulation 1 for $c=0.1$, $0.5$, and $0.9$ as $N$ increases. Errors are averaged over 50 trials, with each point representing the mean error and vertical error bars indicating the standard deviation.
Finally, we examine the accuracy of our method for spike detection. To highlight its effectiveness, we consider one vector ($k=1$) in Algorithm P.3. As before, we consider the three regimes $c=0.1,0.5,0.9$ and vary $N$. For each value of $N$, we generate 50 samples of the covariance matrix $W$ and evaluate both the probability of correctly identifying the number of spikes and the average number of detected spikes. We also demonstrate the efficiency of our approach by comparing its average runtime [^7] The timing tests in this section were done on a MacBook Pro running macOS Sequoia 15.7.1 with 10 cores and 16 GB of RAM with an Apple M4 chip. with the time required to compute eigenvalues across the 50 samples. The results are presented in Figure 5 and Table 1. Our estimator for $\widehat{r}$ exhibits greater accuracy for $c=0.1$ and $c=0.5$ than for $c=0.9$, which can be attributed to the larger gap between the spikes and the density endpoints. Nonetheless, the accuracy improves as $N$ increases and eventually reaches 1 in all cases. Furthermore, even when the probability of correctly estimating the number of spikes is low, the average value of $\widehat{r}$ remains relatively close to the true spike count. These results also highlight the efficiency of the detection algorithm, with its runtime remaining small as $N$ grows, in contrast to the substantial increase in the computational cost of eigenvalue computation.
Figure 5. Left: Spike detection accuracy in Simulation 1 for $c=0.1$, $0.5$, and $0.9$ as NN varies, based on 50 sample realizations with the algorithm run using $k=1$. Right: Comparison of the average computation time for eigenvalue calculation and spike detection across the samples for $c=0.5$.
NN $c=0.1$ $c=0.5$ $c=0.9$
$200$ $3.00$ $(1.00)$ $3.26$ $(0.76)$ $3.20$ $(0.46)$
$2000$ $3.00$ $(1.00)$ $3.06$ $(0.94)$ $3.10$ $(0.92)$
$4000$ $3.00$ $(1.00)$ $3.00$ $(1.00)$ $3.04$ $(0.96)$
$6000$ $3.00$ $(1.00)$ $3.00$ $(1.00)$ $3.02$ $(0.98)$
$8000$ $3.00$ $(1.00)$ $3.00$ $(1.00)$ $3.04$ $(0.96)$
Table 1. Estimated number of spikes for $W$ in Simulation 1 for different values of $c$ and $N$, using a single vector ($k=1$) in the detection algorithm. Each table entry represents the average over 50 samples, with the value in brackets denoting the probability of correctly detecting the true number of spikes.
@@
We set $N=8000$ and $M=16000$, with $X\in\mathbb{R}^{N\times M}$ consisting of iid standard normal entries  [^8] Similar results were obtained when the entries of $X$ followed a Rademacher distribution or a Beta distribution with shape parameters $\alpha=\beta=1/2$.. The parameter $\delta$ ranges from $1.5$ to $3$, crossing the BBP transition, where the number of spikes increases from $2$ to $3$ at $\delta=1+\sqrt{c}$ with $c=N/M$. We evaluate the performance of Algorithm P.3 with $k=1,50,100,200$, alongside BEMA0 and DDPA, in estimating the number of spikes across $50$ realizations of the sample covariance matrix. In addition, we examine the average number of detected spikes and compare the computational efficiency of each method based on their average runtime. The results are summarized in Figures 6, 7 and Table 2.
Figure 6. Spike detection accuracy in Simulation 2, averaged over 50 trials for each value of $\delta$ as it varies from 1 to 3. During this process, the sample covariance matrix undergoes the BBP transition, increasing the number of spikes from 2 to 3. We compare the performance of Algorithm P.3 with different numbers of vectors against the BEMA0 and DDPA methods.
Figure 7. Comparison of the average computation time in Simulation 2, measured over 50 trials, for Algorithm P.3 with varying numbers of vectors, alongside BEMA0 and DDPA. The runtimes of BEMA0 and DDPA are nearly identical, as both are primarily determined by the cost of eigenvalue computation.
@@
$\delta$ Lanczos . with . $k=1$ Lanczos . with . $k=50$ Lanczos with . $k=100$ Lanczos with $k=200$ BEMA0 DDPA
$1.50$ $2.00$ $(1.00)$ $2.00$ $(1.00)$ $2.00$ $(1.00)$ $2.00$ $(1.00)$ $2.10$ $(0.90)$ $2.00$ $(1.00)$
$1.75$ $2.00$ $(0.00)$ $2.00$ $(0.00)$ $2.00$ $(0.00)$ $2.00$ $(0.00)$ $2.64$ $(0.64)$ $2.00$ $(0.00)$
$1.90$ $2.00$ $(0.00)$ $2.00$ $(0.00)$ $2.00$ $(0.00)$ $2.00$ $(0.00)$ $3.02$ $(0.98)$ $2.00$ $(0.00)$
$2.00$ $2.06$ $(0.06)$ $2.00$ $(0.00)$ $2.04$ $(0.04)$ $2.04$ $(0.04)$ $3.02$ $(0.98)$ $2.00$ $(0.00)$
$2.10$ $2.72$ $(0.72)$ $2.78$ $(0.78)$ $2.76$ $(0.76)$ $2.78$ $(0.78)$ $3.08$ $(0.92)$ $2.00$ $(0.00)$
$2.25$ $2.98$ $(0.98)$ $2.94$ $(0.94)$ $2.96$ $(0.96)$ $3.00$ $(1.00)$ $3.00$ $(1.00)$ $2.00$ $(0.00)$
$2.50$ $3.00$ $(1.00)$ $3.00$ $(1.00)$ $3.00$ $(1.00)$ $3.00$ $(1.00)$ $3.12$ $(0.88)$ $2.00$ $(0.00)$
$2.75$ $3.00$ $(1.00)$ $3.00$ $(1.00)$ $3.00$ $(1.00)$ $3.00$ $(1.00)$ $3.04$ $(0.96)$ $3.00$ $(1.00)$
Table 2. Estimated number of spikes for $W$ in Simulation 2 as $\delta$ crosses the BBP transition, increasing the number of spikes from $2$ to $3$. We compare Algorithm P.3 with different numbers of vectors against BEMA0 and DDPA. Each table entry represents the average over 50 samples, with the value in brackets indicating the probability of correctly identifying the true number of spikes.
@@
Figure 8. Left: The ESD of $\Sigma$, where $\Sigma$ is an $N\times N$ diagonal matrix with $N=3000$ and entries given by the quantiles of the density defined in (7.5). The first two diagonal entries are modified to 7 and 6 forming the spikes of $\Sigma$. Right: The ESD of $W$ defined in Simulation 3. The ESD is compared against the estimated outliers, their locations, and the approximate ASD obtained using Algorithms P.2 and P.3 with $k=100$.
We consider the sample covariance matrix $W$ under the spiked covariance model in (1.1), where $N=3000$, $M=30000$ and $\sqrt{M}X\in\mathbb{R}^{N\times M}$ has iid standard Gaussian entries. The covariance matrix $\Sigma\in\mathbb{R}^{N\times N}$ is diagonal, and its eigenvalues follow a deterministic distribution consisting of a bulk component and additional spikes. The bulk density is expressed as
@@
where $K$ is the normalization constant, and the spikes of $\Sigma$ are located at $7,\delta$. See Figure 8 for a visualization of the ESD of $\Sigma$ and the ASD of $W$ when $\delta=6$. We approximate the ASD and detect the outliers of $W$ using Algorithms P.2 and P.3 with $k=100$ vectors. Our ASD estimate closely matches the empirical distribution, accurately identifying both the location and number of spikes.
Next, we vary the parameter $\delta=5,7,9,21$ and evaluate the performance of our detection algorithm with a single vector ($k=1$), comparing it to other spike detection methods. For each value of $\delta$, we perform the evaluation over 50 samples. Specifically, we examine the accuracy, the average number of spikes detected, and the runtime for each method. The results are summarized in Table 3.
Our algorithm outperforms the other methods, as it consistently captures the correct number of spikes with a probability close to 1. DDPA performs second best, with its performance improving as $\delta$ increases, reaching a probability of 1 when $\delta=9,21$. This demonstrates the versatility of our approach, as it can handle a wide range of covariance matrices without assuming any specific distribution or structure for $\Sigma$. Additionally, we observe that Algorithm P.3 achieves the shortest runtime, followed by BEMA0 and DDPA, which have similar runtimes as their costs are mainly dominated by the eigenvalue computations. BEMA and Pass&Yao are more computationally demanding, as they rely on Monte Carlo simulations with large matrices.
@@
Probability $0.00$ $0.00$ $0.50$ $0.00$ $0.88$
Average $553.66$ $0.00$ $2.86$ $1.00$ $1.88$
Time (sec) $0.794847$ $288.6114$ $227.1758$ $0.794932$ $0.035953$
@@
Probability $0.00$ $0.00$ $0.36$ $0.76$ $1.00$
Average $553.68$ $0.00$ $3.18$ $1.52$ $2.00$
Time (sec) $0.797435$ $288.3089$ $245.7479$ $0.797535$ $0.033605$
@@
Probability $0.00$ $0.00$ $0.40$ $1.00$ $0.98$
Average $553.74$ $0.00$ $3.04$ $2.00$ $2.02$
Time (sec) $0.784069$ $288.4510$ $238.0594$ $0.784192$ $0.032017$
@@
Probability $0.00$ $0.00$ $0.34$ $1.00$ $1.00$
Average $553.74$ $1.00$ $3.16$ $2.00$ $2.00$
Time (sec) $0.779923$ $290.1505$ $246.6577$ $0.779997$ $0.028166$
Table 3. Probability of correct estimation, average number of spikes detected, and runtime for different methods in Simulation 3, evaluated over 50 samples for each value of $\delta$.